# MCTS-SQL: An Effective Framework for Text-to-SQL with Monte Carlo Tree Search

### v1

@@ 4.3 Generator
Specifically, the Direct Generator employs chain-of-thought (COT) prompting, which is proven effective during SQL generation. The Selector provides a selected subset of tables and columns to minimize confusing information during the generation. And the user’s query is added to the same prompt. The LLM processes this input to generate SQL queries, accompanied by a detailed rationale.
Additionally, we employ a few-shot learning strategy, using several in-context examples to improve the LLM’s understanding of task-specific instructions and enhance its generalization capabilities.
4.3.2 B. MCTS-Refiner
The MCTS-Refiner aims to refine SQL queries using execution errors, and the self-critique mechanism is adopted to optimize the query iteratively. The main workflow of the proposed method consists of several stages, detailed as follows:
@@
Backpropagation: The value $r$ of the refined SQL query is backpropagated through the search tree, updating the value information of the parent node and other relevant nodes. If the $P$-value of any child node is changed, the corresponding $P$-value of its parent node is recalculated accordingly. The process can be described as follows:
$P^{\prime}(a)=\frac{1}{2}\left(P(a)+\max_{i\in\text{Children}(a)}P(i)\right)$ (8)
UCT update: After updating the $P$ values for all nodes, we choose the UCT function to measure the combined value of each node, which is used as an important basis for expansion in the next selection stage. The UCT value of a node $a$ is formally defined as:
@@
The algorithm proceeds through all these steps iteratively until the maximum rollout numbers are reached. And the SQL querie with the highest score $r$ is chosen as the final output.
@@ 5 Experiments
To evaluate the performance of our MCTS-SQL, we present the implementation details, explain the experiments performed, and offer a thorough analysis of the results.
5.1 Datasets
We evaluate our MCTS-SQL framework using two NL2SQL benchmarks: Spider and BIRD. The Spider contains 7000 training question-query pairs and 1038 development pairs, which are from 200 different databases across 138 domains. In this study, we only concentrate on the development set. In contrast, BIRD includes 95 large-scale databases with high-quality Text-to-SQL pairs, covering 37 specialized domains. Unlike Spider, BIRD focuses on real-world database content and provides external knowledge reasoning to bridge natural language queries and database contextual information.
5.2 Evaluation Metrics
To evaluate our proposed method’s performance, we use two metrics: Execution Accuracy(EX) and Valid Efficiency Score(VES). The Execution Accuracy(EX) calculates the percentage of queries where the predicted SQL queries match the correct SQL queries when executed. Valid Efficiency Score(VES) measures the percentage of predicted SQL queries that output sets consisting of the results from the ground-truth SQL queries.
5.3 Base Models
In this paper, we utilized the latest models, GPT-4o-mini and GPt-4o, as the base models for our conducted experiments. Most of our ablation studies experiments are carried out with GPT-4o-mini, which offers a nearly 30% cost reduction compared to GPT-4o. Despite this, GPT-4o-mini proved to be more effective, offering a powerful balance between performance and cost. In future works, we will explore other models and fine-tune an open-source LLM.
5.4 Hyper-parameters
In order to ensure the stability of our experiment results, we standardized the hyper-parameters as follows. The temperature is fixed at 0.0, the top-p parameter is set to 1.0, and the max-token length is 32168. As for the hyper-parameters in the MCTS-Refiners, the child nodes of a node are set to 2, and the max-rollout numbers are 5.
5.5 Results
5.5.1 A. BIRD Results
In Table 1, we present the comparison of our method’s performance on the BIRD against existing approaches. Our proposed MCTS-SQL achieves the highest performance with 69.40% for dev EX and 66.24% for dev VES, demonstrating its superiority over other methods. Even when using the more cost-effective model GPT-4o-mini, our method still performs well, with scores of 63.15% and 60.78%. This highlights the effectiveness of our approach, which can achieve competitive results even with a lighter model.
The detailed performance of different complexity levels is shown in Table 2. The results show a notable improvement. Based on the GPT-4o, the EX of the simple subset is 74.32%, the moderate subset is 65.17%, and a challenging subset is 51.48%. Analyzing the results, we can find a great outperforming of challenging sets, mainly due to the MCTS-based refiner, which is an important aspect of handling real-world applications.
Method dev EX text EX dev VES
@@ 5.5 Results
Table 1: Comparison with the results of existing methods on BIRD of the Execution accuracy and Valid Efficiency Score.
@@
E-SQL+GPT-4o-mini 67.44 56.94 40.00 59.81
E-SQL+GPT-4o 73.02 64.14 48.07 66.29
@@
Table 2: Execution accuracy in BIRD development set.
@@
Table 3: Execution accuracy on both dev and test set of spider
5.5.2 B. Spider Results
Table 3 shows the performance comparison on the Spider dataset. Our MCTS-SQL achieves excellent results with 88.71% on the development set and 86.63% on the test set. We can see that existing work has achieved outstanding performance; even in this situation, our work still performs remarkably well.
@@ 5.6 Ablation study
We conducted an ablation study to analyze the impact of three of our proposed components-Schema, Selector, and MCTS-Refiner. The results, shown in Table 4, highlight the importance of each module in the MCTS-SQL. To reduce costs, we use GPT-4o-mini as the base model. Specifically, removing the carefully designed Schema and using a normal DDL slightly reduces performance. The Selector is more important, especially on both moderate and challenging subsets. The most significant decrease occurs when the MCTS-Refiner is removed, demonstrating its critical role in solving challenging conditions. All these tricks boost the system’s effectiveness in the NL2SQL task and can be transferred to other relevant problems.
@@
Ours + GPT-4o-mini 68.56 57.76 45.83 63.15
w/o schema 66.98 56.27 43.82 61.56($\downarrow$)
w/o Selector 66.71 55.40 40.18 60.79($\downarrow$)
w/o MCTS-Refiner 64.46 52.00 35.44 57.96($\downarrow$)
Table 4: Ablation study using GPT-4o-mini with EX on the development set.
Pipeline Simp. Mod. Chall. All
Ours + GPT-4o-mini 68.56 57.76 45.83 63.15
max-rollouts-5 68.56 57.76 45.83 63.15
max-rollouts-6 68.14 57.81 44.91 62.83
max-rollouts-7 68.42 57.12 45.93 62.88
child nodes-2 68.56 57.76 45.83 63.15
chile nodes-3 68.62 57.57 45.83 63.13
@@
Table 5: Ablation study of hyper-parameters in MCTS with EX on the development set.
In this paper, we use the Monte Carlo Tree Search to enhance the model’s performance, and the number of child nodes and max-rollouts are two hyper-parameters. Table 5 shows the specific results of the ablation study. We find that variations of these settings have minimal effect on the overall performance. For instance, the performance across different max-rollouts and child nodes shows a few changes. Therefore, to lower token consumption, we choose the most efficient settings.
Figure 5: The comparison of normal COT and our MCTS-SQL in solving the same query. The analysis shows that the COT struggle with detailed information and complex constrains due to LLM’s hallucinations
6 Discussion
Figure 5 shows an example of normal COT and MCTS-SQL to solve the same user’s query. Through the error analysis, we can find that the syntax and structure generated by the COT are generally correct, but due to the hallucination of LLM, it struggles with detailed information and complex constraints. The feedback and iterative optimization mechanism of MCTS helps reduce errors caused by these details, thus improving the performance of SQL generation.
7 Conclusion
In conclusion, this paper introduces MCTS-SQL, a novel framework that utilizes the Monte Carlo Tree Search to enhance the LLMs to address challenges in Text-to-SQL tasks. Experimental results from the Spider and BIRD demonstrate the significant advantages of our method, achieving state-of-the-art performance, especially in challenging queries. Specifically,we achieve an EX 69.40% on the BIRD with 51.48% on the challenging set, outperforming existing method by 3.41%. These results show the potential of MCTS-SQL for real-world database applications. However, the MCTS uses multiple iterations which inevitable increases token consumption and time. In the future work, we will explore some methods to solve the constrain.

### v3

@@ 3.3 Direct Generator
Specifically, the Direct Generator employs chain-of-thought prompting. (COT) We assemble the relevant table and field information obtained from the Selector mentioned above with the user input. The LLM processes this input to generate SQL queries, accompanied by a detailed rationale. Additionally, we employ a few-shot learning strategy, using several in-context examples to improve the LLM’s understanding of task-specific instructions and enhance its generalization capabilities.
3.4 MCTS-Refiner
Typically, SQL queries generated by the Direct Generator fail to meet task requirements due to syntactic errors or mismatch with the user’s intent. The MCTS-Refiner aims to refine SQL queries using the self-critique mechanism to optimize the query iteratively. The MCTS-Refiner is conditionally activated via two checks: (1) Execution Failure: If the generated SQL fails to run. (2) Semantic Error: If execution succeeds but an LLM-based Verifier identifies user’s intent mismatch.The main workflow of the proposed method consists of several stages, detailed as follows:
@@ 3.4 MCTS-Refiner
Backpropagation: The value $r$ of the refined SQL query is back-propagated through the search tree, updating the value information of the parent node and other relevant nodes. If the $P$-value of any child node is changed, the corresponding $P$-value of its parent node is recalculated accordingly. The process can be described as follows:
$P^{\prime}(a)=\frac{1}{2}\left(P(a)+\max_{i\in\text{a}.children}P(i)\right)$ (8)
UCT update: Following the existing practice (MCTSr), after updating the $P$ values for all nodes, we choose the UCT function to measure the combined value of each node, which is used as an important basis for expansion in the next selection stage. The UCT value of a node $a$ is formally defined as:
@@
The algorithm proceeds through all these steps iteratively until the maximum rollout numbers are reached. And the SQL queries with the highest score $r$ is chosen as the final output.
3.5 Prefix-Cache
Rethinking the MCTS iterations, we observed that a large part of input prompts remain unchanged, such as database schema, field descriptions, and few-shot examples. Re-computing these repeated components introduces unnecessary latency and computational overhead.
To address this issue, we design a prefix-caching mechanism (as can be seen in Fig.4) that reuses the intermediate K/V states of the transformer decoder layers:
  • First Pass: When the model processes an input prefix for the first time, we cache the K/V states generated for decoder layer.
  • Subsequent Passes: If the same prefix appears again, the cached K/V states are restored to skip redundant computation and focus on the new suffix.
In practice, we define the prefix as the invariant components(e.g.,field descriptions, few-shot examples), while the suffix consists of the parts that change during iterations, such as instructions, evaluations and execution feedback. To align with the prefix-caching mechanism, the overall prompt is structured as: dataset schema + few-shot demonstrations + specific instructions + feedback.
Figure 4: Illustration of the proposed optimization strategies. (a) MCTS-based iterative refinement for SQL generation. (b) Prefix-caching mechanism that reuses cached K/V states for invariant prompt components, reducing redundant computation and improving efficiency during multiple iterations.
@@ 4 Experiments
To evaluate the performance of our MCTS-SQL, we present the implementation details, explain the experiments performed, and offer a thorough analysis of the results. We evaluate our MCTS-SQL frameWORK using two Text-to-SQL benchmarks: Spider and BIRD.
Method dev EX test EX dev VES
@@
Alpha-SQL 69.70 - -
XiYan-SQL 73.34 75.63 -
Chase-SQL 74.46 74.79 -
DSAIR-SQL 74.32 74.12 -
Ours+Qwen-1.5B 40.69 43.72 44.87
Ours+Qwen-3B 46.71 48.37 48.19
Ours+Qwen-7B 53.61 51.79 52.21
@@
Ours+Gemini2.5-pro 72.91 74.74 72.66
Table 1: Comparison with the results of existing methods on BIRD of the Execution accuracy and Valid Efficiency Score.The Qwen models in this table are Qwen-Coder-Instruct.
4.1 Evaluation Metrics
To evaluate our proposed method’s performance, we use two metrics: Execution Accuracy(EX) and Valid Efficiency Score(VES).(BIRD; TESTSUIES) The Execution Accuracy(EX) calculates the percentage of queries where the predicted SQL queries match the correct SQL queries when executed. Valid Efficiency Score(VES) measures the percentage of predicted SQL queries that output sets consisting of the results from the ground-truth SQL queries.
4.2 Base Models
In this paper, we adopt Qwen2.5-Coder-Instruct series as our base models, given their leading performance in the filed of code generation. We evaluate the effectiveness of our framework across multiple model sizes, including 1.5B, 3B and 14B, to demonstrate its ability to enhance lightweight models. To further explore the boundary of our method, we also conduct experiment using more powerful closed APIs, including GPT-4o-mini, GPT-4o and Gemini2.5. In future work, we aim to distill the reasoning precess into a lightweight model, which currently implemented through MCTS refinement.
4.3 Hyper-parameters
In order to ensure the stability of our experiment results, we standardized the hyper-parameters as follows. The temperature is fixed at 0.1, the top-p parameter is set to 1.0, and the max-token length is 32168. As for the hyper-parameters in the MCTS-Refiners, the child nodes of a node are set to 2, and the max-rollout numbers are 5.
4.4 Experimental Results
A. BIRD Results
Table 1 presents a performance comparison of our method in the BIRD dataset against existing approaches. When using lightweight models(1.5B and 3B), MCTS-SQL outperforms ChatGPT-3.5 and even rivals some earlier methods based on GPT-4. This demonstrates that our method can be deployed on resource-constrained edge devices, without relying on large-scale models or costly APIs. Furthermore, when combined with the most powerful Gemini2.5-pro, MCTS-SQL reaching 72.91% execution accuracy (EX) and 72.66% value execution score (VES) on the development set, confirming its superiority over existing methods and its practical utility.
Table 2 shows the detailed performance across different complexity levels. Compared to the baseline, our method achieve significant improvements. Analyzing the results, we observe that the model improves more on simpler examples. This may because mistakes in these cases are mostly about syntax, and the MCTS-Refiner can fix them easily using its feedback-based editing process. However, with a stronger base model, MCTS can still bring clear improvements on harder examples by exploring different ways to fix the errors.
@@ A. BIRD Results
Qwen-1.5B 15.36 14.96 9.78 14.71
Qwen-3B 19.24 16.18 12.49 17.68
Ours + Qwen-1.5B 46.36 34.96 22.78 40.69
Ours + Qwen-3B 53.74 39.62 24.55 46.71
Ours + Qwen-7B 62.98 42.21 30.28 53.61
@@
Ours+Gemini2.5-pro 76.98 69.82 56.84 72.91
Table 2: Execution accuracy in BIRD development set. The Qwen models in this table are Qwen-Coder-Instruct.
B. Spider Results
Table 3 presents the performance comparison on the Spider dataset. When using lightweight models, our method achieves results that are already practically usable. Furthermore, when equipped with GPT-4o as the base model, MCTS-SQL achieves outstanding performance, reaching 89.17% on the development set and 88.74% on the test set. While existing approaches have already demonstrated strong results on this benchmark, our method continues to deliver highly competitive performance.
@@ B. Spider Results
Ours+Gemeni2.5-pro 89.17 88.74
Table 3: Execution accuracy on both dev and test set of spider. The Qwen models in this table are Qwen-Coder-Instruct.
Setting Avg Latency Tokens Ex Accuracy
single-shot 0.63s 197 14.71
no-cache 6.12s 2274 40.69
with-cache 2.84s 864 40.42
Table 4: Impact of Prefix-Cache on Latency, Token Usage, and Execution Accuracy
4.5 Prefix-Cache Effectiveness Evaluation
To evaluate the effectiveness of the proposed prefix-cache mechanism, we conducted controlled experiments focusing on three key aspects: (I) inference latency, (II) token computation reduction, and (III) execution accuracy impact. All experiments were performed on the BIRD . For a fair comparison, we used the same base model (Qwen2.5-Coder-Instruct-1.5B) under two settings: with prefix-cache and without prefix-cache. The detailed results can be seen in Table 4.
The prefix-cache stores the intermediate key-value states for invariant parts of the input, such as database schema and few-shot examples. This enables the model to skip redundant computations during subsequent decoding steps. We measured inference time on 500 multi-turn SQL generation tasks.
This pre-computation, though highly efficient, means the model’s sampling process begins from a fixed distribution. We think that may occasionally guide the model toward a sub-optimal SQL generation path, especially when integrating new, dynamic feedback from the suffix. However, we consider the minor performance is an acceptable trade-off.
@@ 4.6 Ablation Study
To evaluate the role of each component in MCTS-SQL, we perform an ablation study using Qwen-1.5B as the base model (Table 5). The results confirm that all modules are essential and reflect a core design principle: progressively reducing the search space to mitigate the limited reasoning capacity of lightweight models.
Removing the Selector lowers overall accuracy from 40.7% to 36.2%, as the model must handle the full schema with redundant tables and columns. This highlights the Selector’s importance in filtering irrelevant schema elements and simplifying prompts.Eliminating the Direct Generator further reduces accuracy to 34.8%, since the refinement process lacks a strong starting point. By providing an initial candidate, the Direct Generator decreases the search depth required by the MCTS-Refiner.The MCTS-Refiner proves most critical: without it, accuracy drops sharply to 16.8%, close to single-shot generation. This validates the necessity of iterative, feedback-driven refinement for correcting syntax and semantic errors.Finally, removing the Prefix-Cache has minimal impact on accuracy (40.7% → 40.4%) but significantly decrease the latency, demonstrating its role in improving efficiency without sacrificing performance.Overall, these findings show that each component contributes to an efficient and accurate Text-to-SQL process, enabling a lightweight model like Qwen-1.5B to deliver competitive performance under resource constraints.
Setting Simp. Mod. Chall. All
Full Pipeline 46.5 35.0 22.8 40.7
w/o Selector 42.1 30.1 18.1 36.2
w/o Direct Generator 39.9 29.8 17.5 34.8
w/o MCTS-Refiner 19.8 13.5 8.2 16.8
w/o Prefix-Cache 46.4 35.0 22.8 40.4
Single-shot 17.2 12.1 7.1 14.7
Table 5: Ablation study on BIRD dev set using Qwen-1.5B as base model.
5 Error Analysis
We conducted an error analysis of the single-prediction model and found that 42% of the errors were caused by syntax mistakes, wrong field selection, or misunderstanding of the schema. MCTS-SQL improves Text-to-SQL performance by using a trial-and-error feedback mechanism to guide the search for better SQL queries. Instead of relying on a single-shot generation process, it explores multiple candidate SQL queries and keeps those that are actually effective based on execution results. This allows the method to recover from some generation errors. The rate of syntax errors and errors in table or database selection has significantly decreased to 13%, demonstrating the importance of our approach in addressing common pattern problems. We observe a change in error types during refinement: while syntax and simple field errors decrease rapidly, remaining failures are dominated by deep semantic or schema-linking issues. Specifically, MCTS is effective at correcting errors such as spelling mistakes, operator errors . However, it still struggles with errors caused by ambiguous natural language queries as well as complex multi-table joins involving foreign keys.
6 Conclusion
In conclusion, we propose MCTS-SQL, a novel framework to enhance Text-to-SQL performance of lightweight models. Our approach employs Monte Carlo Tree Search for iterative SQL refinement with three modules: Selector, Direct Generator, and MCTS-Refiner. To reduce the overhead of multi-step refinement, we design a token-level prefix-cache mechanism, which significantly improves inference efficiency. Experiments on SPIDER and BIRD show that, even with a 1.5B model, MCTS-SQL outperforms ChatGPT-3.5, and achieves competitive results with Gemini 2.5. These findings demonstrate that MCTS-SQL, together with prefix-cache, provides a practical and efficient solution for real-world applications under resource constraints. Overall, MCTS offers an approach that empowers lightweight models with stronger capabilities, providing an excellent baseline for other complex cross-domain research.
The appendix presents some detailed information and experimental conclusions about the proposed MCTS-SQL.
Appendix A Details on experiments
A.1 An example of schema
This appendix provides an example of the proposed semi-structured schema representation used throughout our framework. It demonstrates how table-level and column-level information is organized to translate the database structure clearly to the language model. For each table, we include its name and an optional description, followed by a list of column entries, each consisting of the column name, data type, description, and example values. Furthermore, explicit foreign key mappings are listed to illustrate the relationships between tables, enabling the model to reason over the database schema more effectively.
Figure 5: An example of proposed database schema format.The format consists of table names, descriptions and column level details(name, data type, description, and examples) to represent the hierarchical information of databases.
A.2 Ablation study of hyper-parameters in MCTS
This appendix section presents additional analysis on the sensitivity of two key hyper-parameters in our Monte Carlo Tree Search (MCTS) module: the maximum number of rollouts and the number of child nodes explored at each step. As shown in Table 1, we systematically vary these parameters and evaluate their effects on performance across different query difficulty levels. The results indicate that the overall performance is relatively stable under such changes, with only slight fluctuations observed. This demonstrates the robustness of our method to these hyper-parameter settings.
@@ A.2 Ablation study of hyper-parameters in MCTS
Ours + Qwen2.5-1.5B-coder 46.36 34.96 22.78 40.69
max-rollouts-5 46.36 34.96 22.78 40.69
max-rollouts-6 46.12 35.02 22.51 40.53
max-rollouts-7 46.44 34.61 22.93 40.64
child nodes-2 46.36 34.96 22.78 40.69
child nodes-3 46.53 34.12 22.89 40.54
@@
Table 6: Ablation study of hyper-parameters in MCTS with EX on the development set using Qwen2.5-1.5B-coder.
A.3 An example of MCTS-Refine
To demonstrate the effectiveness of our MCTS-Refiner, we provide an example. Starting from an initial incorrect SQL, the refiner iteratively explores and improves candidate queries by evaluating execution feedback and applying guided modifications. Through this process, it successfully constructs a correct SQL that joins relevant tables, applies appropriate filters, aggregates and sorts the results, and adds a constraint on the number of charter schools. This example highlights how MCTS-Refiner progressively resolves intent understanding, schema linking, and syntactic issues, ultimately producing an executable and accurate SQL query.
Figure 6: The figure illustrates how the MCTS-Refiner incrementally improves SQL generation by exploring candidate queries through a tree search. Starting from an initial faulty query, it applies guided edits based on execution feedback and critique until a correct SQL is formed that satisfies the input question.
A.4 Baselines
In this appendix, we provide a detailed description of the baselines used in the experiments presented in the paper. Baseline models serve as a benchmark for evaluating the effectiveness of our proposed approach, MCTS-SQL. The selection of these baselines is based on their relevance to the Text-to-SQL task and their ability to represent widely recognized methods in the field. Each baseline is briefly introduced, along with its specific configuration and the rationale for its inclusion in our study. The performance of these models is compared against that of our proposed framework to highlight the improvements achieved through MCTS-SQL.
  • DIN-SQL utilizes zero-shot and few-shot prompting with large language models like GPT-4 and CodeX. While effective for basic queries, it often lags behind fine-tuned models on complex tasks and benchmarks like Spider due to its lack of fine-tuning.
  • DAIL-SQL enhances LLM-based Text-to-SQL by systematically comparing existing prompt strategies, optimizing question representation, and employing supervised fine-tuning.
  • MAC-SQL introduces a multi-agent framework with three components: a Decomposer to break down complex queries, a Selector to focus on relevant database parts, and a Refiner to fine-tune the generated queries. This approach excels on more intricate queries by leveraging agent collaboration and fine-tuned models like SQL-Llama.
  • MCS-SQL employs multiple prompts to explore a wider search space, combined with schema linking and multiple-choice selection to filter and optimize generated SQL queries. This approach has set new benchmarks on the BIRD and Spider datasets.
  • SQL-o1 leverages a self-reward-driven MCTS search mechanism to enhance reasoning in Text-to-SQL tasks. It includes multi-step exploration and dynamic pruning for faster inference, improving generalization and surpassing GPT-4 in performance on the BIRD dataset.
  • Alpha-SQL focuses on zero-shot Text-to-SQL using MCTS. It dynamically guides SQL construction actions with LLM-as-Action-Model, while using a self-supervised reward function to refine the SQL generation process.
  • XiYan-SQL uses an ensemble of multiple generators and fine-tuning strategies to improve SQL candidate diversity and quality. The approach integrates in-context learning (ICL) with M-Schema, a semi-structured schema representation, and includes a refiner to correct errors and select the best candidate.
  • CHASE-SQL adopts a multi-agent framework that generates diverse SQL candidates through a divide-and-conquer approach. It incorporates chain-of-thought reasoning, synthetic example generation, and pairwise candidate comparison, resulting in better SQL query generation and selection.
  • DSAIR focuses on metadata extraction techniques to support SQL generation, addressing challenges in database understanding. It explores profiling, query log analysis, and SQL-to-text generation, reducing the need for human-generated metadata and improving SQL generation quality.
While existing baselines have advanced the Text-to-SQL task through prompt engineering, agent collaboration, supervised fine-tuning, and search strategies, they often rely on either large-scale models or carefully designed training pipelines. In contrast, our motivation lies in enabling lightweight models to effectively handle complex SQL generation without large parameter counts or costly training. Unlike methods such as Alpha-SQL and SQL-o1 that use MCTS mainly in zero-shot or reward-driven settings, MCTS-SQL introduces a structured, feedback-guided refinement process tailored for small models, leveraging schema pruning, strong initialization, and efficient search. This design allows MCTS-SQL to strike a unique balance between performance and resource efficiency, distinguishing it from prior work that either emphasizes prompt diversity or model scale.
Appendix B Details of prompts
B.1 Prompt of selector
You are an expert of SQL. Given a detailed schema description, you should finish the task in the [instruction]. [DB_ID]{db_id} [Schema]{desc_str} [Foreignkeys]{fk_str} Here are few examples: {schema describe} [Instruction]
Your responsibilities include:
1. Discard any table schema that is not related to the user question and accompanying evidence. 2. For each relevant table: If the number of columns is 10 or fewer, mark the table as "keep-all". Otherwise, select the top 6 most relevant columns and list them in descending order of relevance. 3. Ensure at least 3 tables are included in the final output. 4. Output must be in strict JSON format. example’s answers are:
{example’s answer}
[Question]
{query}
[Evidence]
{evidence}
[Answer]
B.2 Prompt of direct generator
You are an expert of SQL. Given a detailed schema description, you should finish the task in the [instruction]. [DB_ID]{db_id} [Schema]{desc_str} [Foreignkeys]{fk_str} Here are few examples: {schema describe} [Instruction]
Your responsibilities include:
1.Understand the schema and foreign key relationships to identify relevant tables and columns. 2.Carefully align the user question with the corresponding elements in the schema. 3.Resolve ambiguities based on schema structure and naming conventions. 4.Compose a complete and executable SQL query that correctly answers the user’s question. 5.Ensure the query respects SQL syntax and follows best practices (e.g., proper JOINs, WHERE conditions, GROUP BY, ORDER BY). 6. Do **not** output any explanation or commentary — only the SQL query. 7.If multiple interpretations exist, select the most plausible one based on provided evidence. example’s answers are:
{example’s answer}
[Question]
{query}
[Evidence]
{evidence}
[Answer]
B.3 Prompt of verifier
You are an expert of SQL. Given a detailed schema description, you should finish the task in the [instruction]. [DB_ID] {db_id} [Schema] {desc_str} [Foreignkeys] {fk_str} Here are few examples: schema describe [Instruction]
Your responsibilities include: 1.Understand the schema and table relationships based on the schema description and foreign key mappings.
2.Analyze the user question and determine what the expected result should be. 3.Review the given SQL query and check whether it correctly retrieves the expected result for the question. 4.Verify that the query uses the correct tables, columns, filters, aggregations, joins, groupings, and ordering. 5.Detect any mismatches, hallucinations (e.g., using non-existent columns), logic errors, or inconsistencies with the user intent. 6.If the SQL is correct, answer ‘YES‘. If the SQL is incorrect, answer ‘NO‘ and briefly explain the main error(s). 7.Do **not** rewrite or fix the SQL — only judge its correctness and explain why. example’s answers are:
{example’s answer}
[Question]
{query}
[Evidence]
{evidence}
[SQL]
{sql}
[Answer]
B.4 Prompt of critique in MCTS
You are an expert of SQL. Given a detailed schema description, you should finish the task in the [instruction]. [DB_ID] {db_id} [Schema] {desc_str} [Foreignkeys] {fk_str} [Instruction]
Your responsibilities include: 1. Identify specific parts of the SQL that are responsible for the error.
2. Check if any schema misunderstanding, incorrect column/table usage, invalid expressions, or logical flaws are present. 3. Provide **constructive and detailed** feedback to help improve the SQL, with reference to schema if necessary. 4. Do **not** rewrite or fix the SQL — only explain what’s wrong and what to consider correcting. [Question]
{query}
[Evidence]
{evidence}
[SQL]
{sql}
[SQL_error] sql_error [Answer]
B.5 Prompt of refine in MCTS
You are an expert of SQL. Given a detailed schema description, you should finish the task in the [instruction]. [DB_ID] {db_id} [Schema] {desc_str} [Foreignkeys] {fk_str} [Instruction]
Your responsibilities include:
1.Continue to comply with the requirements in the original instructions. 2. Give a direct and concise SQL to the instructions, do not refer to or discuss the criticisms. 3. Do not repeat the original instruction or answer statements. 4.Response with only the answer. [Question]
{query}
[Evidence]
{evidence}
[SQL]
{sql}
[SQL_error] sql_error [Critique] critique [Answer]
B.6 Prompt of evaluation in MCTS
You are an expert of SQL. Given a detailed schema description, you should finish the task in the [instruction]. [DB_ID] {db_id} [Schema] {desc_str} [Foreignkeys] {fk_str} [Instruction]
Your responsibilities include:
1. Based on executing SQL and the error information, please provide a reward score between -100 and 100 for the anwer quality, using very strict standards. 2.Do not give a full score above 95. Make sure the reward score is an integer." 3.Return *ONLY* the score." [Question]
{query}
[Evidence]
{evidence}
[SQL]
{sql}
[SQL_error] sql_error [Answer]